# DREAM: Scalable Red Teaming for Text-to-Image Generative Systems via Distribution Modeling

### v1

@@ 4.3 Red Team LLM Optimization
where $\lambda$ and $\beta$ are balancing hyperparameters. However, optimizing Eq. (5) is non-trivial. One intuitive approach would be to use backpropagation-based methods to obtain exact gradients and then update the LLM’s parameters. However, applying backpropagation-based optimization directly is challenging. This is because the full red-teaming pipeline includes multiple components, including autoregressive language generation, multi-step diffusion denoising, and energy models. For instance, generating a 10-token prompt with an autoregressive LLM involves 10 sequential decoding steps, each with its own activations. Likewise, synthesizing an image via diffusion models typically requires 30 iterative denoising steps. Storing the full computation graph for just a single forward pass through this pipeline would easily demand thousands of gigabytes GPU memory, making end-to-end backpropagation-based training memory-prohibitive even for small-scale models and small batch sizes. Moreover, certain components like keyword-based safety filters are non-differentiable, further hindering backpropagation.
To enable effective gradient-driven optimization while avoiding the need for backpropagation through the entire pipeline, we propose a novel framework based on Simultaneous Perturbation Stochastic Approximation (SPSA) [30]. SPSA is a classical zero-th order optimization method that allows estimates of high-dimensional gradients using only forward evaluations. However, Vanilla SPSA is validated to suffer from instability and slow convergence in our red-teaming setup, due to the highly stochastic nature of both LLMs and diffusion-based generation (see experiments in Section 5.7). To mitigate this, we propose a simple yet effective variant, GC-SPSA, which incorporates an adaptive sampling schedule as well as a history-aware gradient calibration mechanism to reduce gradient variance while maintaining efficiency. In the following, we first introduce SPSA and analyze its problems, and then we propose our GC-SPSA, and finally provide a theoretical analysis and convergence guarantee to support our design.
Definition 2 (SPSA [30]).
Given an objective function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and parameters $\theta\in\mathbb{R}^{d}$, SPSA uses the following randomized two-point finite-difference approximation to compute the unbiased estimate of gradient $\nabla_{\theta}\mathcal{L}(\theta)$:
@@
where $\epsilon>0$ is a small perturbation magnitude, $\Delta\in\mathbb{R}^{d}$ is a random perturbation vector sampled from a zero-mean Gaussian distribution.
Previous works have proved that SPSA provides an unbiased estimate of the true gradient, i.e., $\mathbb{E}_{\Delta}[g(\theta)]=\nabla_{\theta}\mathcal{L}(\theta)$ [30]. In our setting, SPSA is particularly advantageous: it avoids backpropagation entirely and requires only forward evaluations under perturbed parameters, making it well-suited for bypassing the long and possibly non-differentiable pipeline. Moreover, it eliminates the need to store intermediate forward activations, further reducing memory consumption compared to backpropagation-based methods [70].
@@
To mitigate this problem, we propose a simple yet effective method, GC-SPSA, which stabilizes SPSA with a novel adaptive gradient calibration algorithm. Our key insight is that Eq. (5) primarily steers the LLM toward a subspace of prompts that are likely to elicit harmful images from the target model. Since the pretrained LLM already possesses strong priors, such subspaces are shown to typically reside within a relatively flat and continuous basin in the loss landscape nearby the pre-trained parameters [71, 72, 73]. Therefore, we hypothesis that it is more crucial to ensure the reliability of early optimization steps to accurately locate this subspace, yet latter updates may tolerate more variance and can be calibrated with early reliable gradients.
@@
1:Input: Initial model parameters $\theta_{0}$, initial sampling budget $n_{0}$, learning rate $\eta$, correction strength $\gamma$, smoothing factor $\rho$, decay factor $T_{\text{dec}}$, max steps $T_{\text{max}}$
2:Output: Optimized generator parameters $\theta_{T_{\max}}$
3$\triangleright$ Initialize gradient estimate and effective sample size
4:$\hat{g}_{0}\leftarrow\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}g_{0,i}(\theta_{0})$
5:$w_{0}\leftarrow n_{0}$
6:for $t=1$ to $T_{\text{max}}$ do
7$\triangleright$ Determine sampling budget for current step
8:   $n_{t}\leftarrow\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T}}\right\rfloor\right)$
9$\triangleright$ Estimate gradients via SPSA with $n_{t}$ queries
10:   for $i=1$ to $n_{t}$ do
11:      Sample perturbation $\Delta_{t,i}\sim\text{Gaussian}^{d}$
12:      $\mathcal{L}^{+}_{t,i},\mathcal{L}^{-}_{t,i}\leftarrow\mathcal{L}(\theta_{t}+\epsilon\Delta_{t,i}),\mathcal{L}(\theta_{t}-\epsilon\Delta_{t,i})$
13:      $g_{t,i}\leftarrow\frac{\mathcal{L}^{+}_{t,i}-\mathcal{L}^{-}_{t,i}}{2\epsilon}\cdot\Delta_{t,i}$
14:   end for
15$\triangleright$ Aggregate and calibrate gradients
16:   $\hat{g}_{t}\leftarrow\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}g_{t,i}+\gamma\cdot\frac{w_{t-1}}{w_{t-1}+n_{t}}\cdot\hat{g}_{t-1}$
17:   $w_{t}\leftarrow\rho\cdot w_{t-1}+(1-\rho)\cdot n_{t}$
18$\triangleright$ Update model parameters
19:   $\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\hat{g}_{t}$
20:end for
21:Return $\theta_{T_{\max}}$
Specifically, for the $t$-th update, we first estimate the gradient $n_{t}$ times via Eq. (6), obtaining a set of stochastic gradient estimates $\{g_{t,1}(\theta_{t}),g_{t,2}(\theta_{t}),\dots,g_{t,n_{t}}(\theta_{t})\}$. The number of queries $n_{t}$ is controlled by an exponentially decaying schedule: $n_{t}=\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T_{\text{dec}}}}\right\rfloor\right)$, where $n_{0}$ is the initial number of sampling times and $T_{\text{dec}}$ governs the decay rate. This scheduling allocates a higher sampling budget to early iterations and gradually reduces the number of samples as optimization stabilizes. In our experiments, we find that setting $n_{0}=4$ and $T_{\text{dec}}=10$, i.e., starting with 4 samples and halving the sampling budget every 10 steps, yields stable and efficient optimization performance (see experiments in Sec. 5.7). To further reduce the variance of the later estimated gradient, inspired by confidence-aware optimal Bayesian fusion [74], we introduce a gradient calibration mechanism. Specifically, we treat each new gradient estimate as a noisy observation and combine it with historical information using a confidence-aware correction term:
@@
where $\hat{g}_{t-1}$ is the accumulated gradient estimate from previous iterations ($\hat{g}_{0}=\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}g_{0,i}(\theta_{0})$), $w_{t-1}$ denotes its effective sample size (initialized as $w_{0}=n_{0}$), $\eta$ is the learning rate. The update of $w_{t}$ follows an exponential moving average rule. We use the term $\frac{w_{t-1}}{w_{t-1}+n_{t}}$ to approximate the relative confidence of historical vs. current gradients, and $\gamma$ controls the overall strength of the correction. Intuitively, our confidence-weighted fusion scheme anchors the current noisy gradient estimate towards the historically aggregated direction, especially when the current estimate is based on fewer samples (i.e., lower confidence). As formally analyzed in Theorem 1, the GC-SPSA estimator achieves a strictly higher signal-to-noise ratio (SNR) than the Vanilla SPSA estimator for all $t\geq 1$, which helps mitigate gradient noise and promotes a more consistent optimization path.
@@
Let $\bar{g}_{k}$ be the Vanilla SPSA estimator and $\hat{g}_{k}=\bar{g}_{k}+H_{k}\hat{g}_{k-1}$ be the GC-SPSA estimator, with $\hat{g}_{0}=\bar{g}_{0}$ and $H_{k}>0$. Then for all $t\geq 1$, the SNR difference between the GC-SPSA and the Vanilla-SPSA admits the explicit positive lower bound:
@@
where the weights are defined as $h_{t}=1$ and $h_{k}=\prod_{j=k+1}^{t}H_{j}$. Here, $P_{t}=\sum_{k=0}^{t}h_{k}$ is the cumulative weight sum and $V_{k}=V_{\text{single}}/n_{k}$ is the gradient variance at step $k$.
The detailed proof is in Appendix A.1. Furthermore, we also provide theoretical global convergence guarantee and convergence rate analysis for our proposed GC-SPSA to optimize our objective functions in Eq. (5) under mild assumptions, as formally shown in the following theorem:
@@
Consider an objective $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ for all $\theta\in\mathbb{R}^{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix. Then, GC-SPSA will converge (i.e., $\min_{t\in[T]}\mathbb{E}[\|g(\theta_{t})\|^{2}]\leq\delta.$) after
$T=\Theta\left(\frac{\mathcal{L}(\theta_{1})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}}\right).$ (9)
iterations. Here, $\mathcal{L}^{*}$ is the global minimum, and the constants $\Upsilon>0$ and $\Xi>0$ denote upper bounds on the average squared norm of the gradient estimator and the accumulated noise variance, respectively.
For training, we begin by perturbing the current model parameters $\theta_{t}$ and estimate the gradient $n_{t}$ times using Eq. (6). Since the objective $\mathcal{L}(\theta)$ involves an intractable expectation, we approximate it via Monte Carlo sampling over a batch of sampled prompts (Alg. 3). The resulting gradient estimates are then averaged and calibrated using our confidence-aware update rule in Eq. (LABEL:eq:spsa-gc). Finally, the calibrated gradient is used to update the parameters $\theta_{t}$. We summarize the complete training procedure in Alg. 2.
@@ 4.4 Inference-Time Adaptive Temperature Scaling
After training, the optimized red team LLM is steered to a distribution of the prompts that can induce unsafe outputs. The final step is to utilize this model to generate red teaming prompts. However, in practice, we find that the model may still fail to adequately explore its support during generation. This is due to the absence of awareness across generations: the model essentially does not know what has already been generated. Therefore, it may produce repetitive tokens that has limited marginal benefits.
To further enhance diversity, we propose an inference-time strategy that encourages diversity via adaptive temperature scaling. Recall that in autoregressive decoding, the model iteratively predicts the next token distribution with $p_{\theta^{*}}({x_{t}}=v|x_{<t})=\frac{\exp\left(\boldsymbol{z}_{t}[v]/\tau\right)}{\sum_{j\in\mathcal{V}}\exp\left(\boldsymbol{z}_{t}[j]/\tau\right)}$, where the temperature hyperparameter $\tau_{t}$ controls the sharpness of the token distribution $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$ at decoding step $t$. Intuitively, lower temperatures make the model more confident (peaky), while higher temperatures flatten the distribution to encourage exploration. As such, we maintain a global token frequency vector $\mathbf{f}\in\mathbb{N}^{|\mathcal{V}|}$, tracking the number of times each vocabulary token has appeared across previous generations. At decoding step $t$, given raw logits $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$, we compute the relative frequency of the top-scoring token $v_{t}=\arg\max_{j}\boldsymbol{z}_{t}[j]$, and use it to scale the temperature:
@@
where $\alpha>0$ is a sensitivity coefficient and $\tau_{\min}$ ($0.8$ in this paper) prevents degeneracy. The final logits are adjusted as $\tilde{\boldsymbol{z}}_{t}=\boldsymbol{z}_{t}/\tau_{t}$ before sampling. This penalizes frequent tokens, promoting underexplored generations without modifying training or architecture. Empirically, we find it improves prompt diversity with minimal overhead and little loss in effectiveness.

### v2

@@ 4.3 Red Team LLM Optimization
where $\lambda$ and $\beta$ are balancing hyperparameters. However, optimizing Eq. (5) is non-trivial. One intuitive approach would be to use backpropagation-based methods to obtain exact gradients and then update the LLM’s parameters. However, applying backpropagation-based optimization directly is challenging. This is because the full red-teaming pipeline comprises multiple components such as autoregressive language generation, multi-step diffusion denoising, and energy models, each requiring storage of numerous intermediate activations for backpropagation-based gradient computation. For instance, generating an image via Stable Diffusion v1.5 typically requires 30 denoising steps, each producing high-dimensional feature maps (e.g., $\sim$2.3 GB). Even under a moderate batch size of 32, this already amounts to over 2 TB of GPU memory for a single pass, making end-to-end backpropagation-based training memory-prohibitive. Moreover, certain components like keyword-based safety filters are non-differentiable, further hindering backpropagation.
To enable effective gradient-driven optimization while avoiding the need for backpropagation through the entire pipeline, we propose a novel framework based on Simultaneous Perturbation Stochastic Approximation (SPSA) [spall1992multivariate]. SPSA is a classical zeroth-order optimization method that allows estimates of high-dimensional gradients using only forward evaluations. However, vanilla SPSA has been empirically observed to suffer from instability and slow convergence in our red-teaming setup, due to the highly stochastic nature of both LLMs and diffusion-based generation (see experiments in Section 5.8). To mitigate this, we propose a simple yet effective variant, GC-SPSA, which incorporates an adaptive sampling schedule as well as a history-aware gradient calibration mechanism to reduce gradient variance while maintaining efficiency. In the following, we first introduce SPSA and analyze its limitations, then present our GC-SPSA, and finally provide theoretical analysis and convergence guarantees to support our design.
Definition 2 (SPSA [malladi2023fine]).
Given an objective function $\mathcal{L}:\mathbb{R}^{d}\rightarrow\mathbb{R}$ and parameters $\theta\in\mathbb{R}^{d}$, SPSA uses the following randomized two-point finite-difference approximation to compute an unbiased estimate of the gradient $\nabla_{\theta}\mathcal{L}(\theta)$:
@@
where $\epsilon>0$ is a small perturbation magnitude, and $\Delta\in\mathbb{R}^{d}$ is a random perturbation vector sampled from a zero-mean Gaussian distribution.
Previous works have proved that SPSA provides an unbiased estimate of the true gradient, i.e., $\mathbb{E}_{\Delta}[g(\theta)]=\nabla_{\theta}\mathcal{L}(\theta)$ [spall1992multivariate]. In our setting, SPSA is particularly advantageous: it avoids backpropagation entirely and requires only forward evaluations at perturbed parameters, making it well-suited for bypassing the long and possibly non-differentiable pipeline. Moreover, it eliminates the need to store intermediate forward activations, further reducing memory consumption compared to backpropagation-based methods [malladi2023fine].
@@
To mitigate this problem, we propose a simple yet effective method, GC-SPSA, which stabilizes SPSA with a novel adaptive gradient calibration algorithm. Our key insight is that Eq. (5) primarily steers the LLM toward a subspace of prompts that are likely to elicit harmful images from the target model. Since the pretrained LLM already possesses strong priors, such subspaces are shown to typically reside within a relatively flat and continuous basin in the loss landscape near the pretrained parameters [xu2024random, jain2024mechanistically, zhang2023fine]. Therefore, we hypothesize that it is more crucial to ensure the reliability of early optimization steps to accurately locate this subspace, yet later updates may tolerate more variance and can be calibrated with early reliable gradients.
@@
1:Input: Initial model parameters $\theta_{0}$, initial sampling budget $n_{0}$, learning rate $\eta$, correction strength $\gamma$, smoothing factor $\rho$, decay factor $T_{\text{dec}}$, current step $t$
2:Output: Optimized generator parameters $\theta_{t}$
[…]
5:while loss not converged do
[…]
7:   $n_{t}\leftarrow\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T_{\text{dec}}}}\right\rfloor\right)$
[…]
9:   for $i=1$ to $n_{t}$ do
10:   Sample perturbation $\Delta_{t,i}\sim\text{Gaussian}^{d}$
11:   $\mathcal{L}^{+}_{t,i},\mathcal{L}^{-}_{t,i}\leftarrow\mathcal{L}(\theta_{t}+\epsilon\Delta_{t,i}),\mathcal{L}(\theta_{t}-\epsilon\Delta_{t,i})$
12:   $g_{t,i}\leftarrow\frac{\mathcal{L}^{+}_{t,i}-\mathcal{L}^{-}_{t,i}}{2\epsilon}\cdot\Delta_{t,i}$
13:   end for
[…]
15:   $\hat{g}_{t}\leftarrow\frac{1}{n_{t}}\sum_{i=1}^{n_{t}}g_{t,i}+\gamma\cdot\frac{w_{t-1}}{w_{t-1}+n_{t}}\cdot\hat{g}_{t-1}$
16:   $w_{t}\leftarrow\rho\cdot w_{t-1}+(1-\rho)\cdot n_{t}$
[…]
18:   $\theta_{t+1}\leftarrow\theta_{t}-\eta\cdot\hat{g}_{t}$
19:   $t\leftarrow t+1$
20:end while
21:Return $\theta_{t}$
Specifically, for the $t$-th update, we first estimate the gradient $n_{t}$ times using Eq. (6), obtaining a set of stochastic gradient estimates $\{g_{t,1}(\theta_{t}),g_{t,2}(\theta_{t}),\dots,g_{t,n_{t}}(\theta_{t})\}$. The number of queries $n_{t}$ is controlled by an exponentially decaying schedule: $n_{t}=\max\left(1,\left\lfloor\frac{n_{0}}{2^{t/T_{\text{dec}}}}\right\rfloor\right)$, where $n_{0}$ is the initial number of sampling times and $T_{\text{dec}}$ governs the decay rate. This scheduling allocates a higher sampling budget to early iterations and gradually reduces the number of samples as optimization stabilizes. In our experiments, we find that setting $n_{0}=4$ and $T_{\text{dec}}=10$, i.e., starting with 4 samples and halving the sampling budget every 10 steps, yields stable and efficient optimization performance (see experiments in Sec. 5.8). To further reduce the variance of the later estimated gradient, inspired by confidence-aware optimal Bayesian fusion [griebel2020kalman], we introduce a gradient calibration mechanism. Specifically, we treat each new gradient estimate as a noisy observation and combine it with historical information using a confidence-aware correction term:
@@
where $\hat{g}_{t-1}$ is the accumulated gradient estimate from previous iterations ($\hat{g}_{0}=\frac{1}{n_{0}}\sum_{i=1}^{n_{0}}g_{0,i}(\theta_{0})$), $w_{t-1}$ denotes its effective sample size (initialized as $w_{0}=n_{0}$), $\eta$ is the learning rate. The update of $w_{t}$ follows an exponential moving average rule. We use the term ${w_{t-1}}/{w_{t-1}+n_{t}}$ to approximate the relative confidence of historical vs. current gradients, and $\gamma$ controls the overall strength of the correction. Intuitively, our confidence-weighted fusion scheme anchors the current noisy gradient estimate towards the historically aggregated direction, especially when the current estimate is based on fewer samples (i.e., lower confidence). As formally analyzed in Theorem 1, the GC-SPSA estimator achieves a strictly higher signal-to-noise ratio (SNR) than the vanilla SPSA estimator for all $t\geq 1$, which helps mitigate gradient noise and promotes a more consistent optimization path.
@@
Let $\|g_{\text{true}}\|$ be the ground-truth gradient, $\bar{g}_{k}$ be the vanilla SPSA estimator, and $\hat{g}_{k}=\bar{g}_{k}+H_{k}\hat{g}_{k-1}$ be the GC-SPSA estimator, with $\hat{g}_{0}=\bar{g}_{0}$ and $H_{k}>0$. Then for all $t\geq 1$, the SNR difference between the GC-SPSA and the vanilla SPSA admits the explicit positive lower bound $\mathcal{D}_{t}$ as follows:
@@
where $h_{t}=1$, $h_{k}=\prod_{j=k+1}^{t}H_{j}$ and $H_{j}=\gamma\frac{w_{j-1}}{w_{j-1}+n_{j}}$. $V_{\rm{single}}$ is the gradient estimation variance of vanilla SPSA, $P_{t}=\sum_{k=0}^{t}h_{k}$ is the cumulative weight sum, and $V_{k}=V_{\rm{single}}/n_{k}$ is the gradient variance of GC-SPSA at step $k$.
The detailed proof is in Appendix A.1. Furthermore, we also provide a theoretical guarantee of global convergence and convergence rate analysis for GC-SPSA under mild assumptions, as formally shown in the following theorem:
@@
Consider an objective $\mathcal{L}:\mathbb{R}^{d}\to\mathbb{R}$ satisfying $\nabla^{2}\mathcal{L}(\theta)\preceq\ell I_{d}$ for all $\theta\in\mathbb{R}^{d}$, where $I_{d}$ denotes the $d$-dimensional identity matrix. Then, for a stationarity level $\delta$, GC-SPSA will converge (i.e., $\min_{t\in[T]}\mathbb{E}[\|g(\theta_{t})\|^{2}]\leq\delta$) after
$T=\Theta\left(\frac{\mathcal{L}(\theta_{0})-\mathcal{L}^{*}}{\eta\,\zeta_{\min}\,\delta-C_{\max}\,\Upsilon-\frac{\ell\,\eta^{2}\,\Xi}{2}}\right).$ (9)
iterations. Here, $\mathcal{L}^{*}$ denotes the global minimum, $\zeta_{\min}$ is the minimum descent coefficient, $C_{\max}$ bounds the cross-term coefficient, and $\Upsilon$ and $\Xi$ denote upper bounds on the average squared gradient-estimator norm and accumulated noise variance, respectively.
For each training step, we begin by perturbing the current model parameters $\theta_{t}$ and estimate the gradient $n_{t}$ times using Eq. (6). Since the objective $\mathcal{L}(\theta)$ involves an intractable expectation, we approximate it via Monte Carlo sampling over a batch of sampled prompts (see more details in Alg. 3). The resulting gradient estimates are then averaged and calibrated using our confidence-aware update rule in Eq. (LABEL:eq:spsa-gc). Finally, the calibrated gradient is used to update the parameters $\theta_{t}$. We summarize the complete training procedure in Alg. 2.
@@ 4.4 Inference-Time Adaptive Temperature Scaling
After training, the red-team LLM is steered toward the desired prompt distribution. The final step is sampling from this model to generate red-teaming prompts. However, we empirically find that the model may still fail to adequately explore its support during generation. This is because each sample is produced independently and without awareness of previously generated prompts, leading to repetitive or redundant tokens that have limited marginal benefits.
To further enhance diversity, we propose an inference-time strategy that encourages diversity via adaptive temperature scaling. Recall that in autoregressive decoding, the model iteratively predicts the next token distribution as $p_{\theta^{*}}({x_{t}}\!=\!v|x_{<t})=\frac{\exp\left(\boldsymbol{z}_{t}[v]/\tau_{t}\right)}{\sum_{j\in\mathcal{V}}\exp\left(\boldsymbol{z}_{t}[j]/\tau_{t}\right)}$, where the temperature hyperparameter $\tau_{t}$ controls the sharpness of the token distribution $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$ at decoding step $t$. Intuitively, lower temperatures make the model more confident (peaky), while higher temperatures flatten the distribution to encourage exploration. As such, we maintain a global token frequency vector $\mathbf{f}\in\mathbb{N}^{|\mathcal{V}|}$, tracking the number of times each vocabulary token has appeared across previous generations. At decoding step $t$, given raw logits $\boldsymbol{z}_{t}\in\mathbb{R}^{|\mathcal{V}|}$, we compute the relative frequency of the top-scoring token $v_{t}=\arg\max_{j}\boldsymbol{z}_{t}[j]$, and use it to scale the temperature:
@@
where $\alpha>0$ is a sensitivity coefficient and $\tau_{\min}$ prevents degeneracy. The final logits are adjusted as $\tilde{\boldsymbol{z}}_{t}=\boldsymbol{z}_{t}/\tau_{t}$ before sampling. This penalizes frequent tokens, promoting underexplored generations without modifying training or architecture. Empirically, we find it improves prompt diversity with minimal overhead and little loss in effectiveness.